OpenStudy (anonymous):

The diagonals of a rhombus are 6 inches and 8 inches.What is the perimeter of the rhombus?

jimthompson5910 (jim_thompson5910):

the diagonals are perpendicular bisectors of each other so they meet at a 90 degree angle and they cut each other in half 

jimthompson5910 (jim_thompson5910):

you can then use the pythagorean theorem to find x a^2 + b^2 = c^2 3^2 + 4^2 = x^2 I'll let you finish

OpenStudy (anonymous):

what do u mean do i multiply it

jimthompson5910 (jim_thompson5910):

what is 3^2 ?

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

3^2 = 3*3 = ??

OpenStudy (anonymous):

9

OpenStudy (anonymous):

what do u mean by 3^2

jimthompson5910 (jim_thompson5910):

4^2 = 4*4 = ??

jimthompson5910 (jim_thompson5910):

3^2 = 3 squared

OpenStudy (anonymous):

8

OpenStudy (anonymous):

16

jimthompson5910 (jim_thompson5910):

4*4 = ???

OpenStudy (anonymous):

4 squared

jimthompson5910 (jim_thompson5910):

3^2 + 4^2 = x^2 9 + 16 = x^2 25 = x^2 x^2 = 25 x = ???

OpenStudy (anonymous):

x=5

jimthompson5910 (jim_thompson5910):

yep, so each side is 5